OpenStudy (anonymous):

Find the slope for this equation: 3x - 5y = 30

OpenStudy (texaschic101):

In y = mx + b form, the slope will be in the m position. 3x - 5y = 30 -- add 5y to both sides 3x = 5y + 30 -- subtract 30 from both sides 3x - 30 = 5y -- now divide both sides by 5 3/5x - 30/5 = y -- rearrange y = (3/5)x - 6 so what number is in the m position ?

OpenStudy (texaschic101):

y = mx + b y = (3/5)x - 6 the number in the m position is your slope do you know what that number is ?

OpenStudy (anonymous):

Is it 3?